OpenStudy (anonymous):

calculate area under (ln(x))²/x, between x=1, x=e

OpenStudy (anonymous):

\[\Large A=\int_1^e \frac{(\ln(x))^2}{x}dx \]

OpenStudy (anonymous):

use an appropriate u-substitution. can you think of one?

OpenStudy (reemii):

you can use the integration-by-part formula, twice. The first time, take \(f(x)=(\ln(x))^2\) and \(g'(x)=\frac 1x\). Then \[A=[\ln(x)^2\ln(x)]^e_1 - \int_1^e (2\ln(x)\frac 1x)(-\frac 1{x^2})dx \] put again the \(\ln(x)\) in \(f\) and \(g'\) is the rest.

OpenStudy (anonymous):

I believe \[\Large u=\ln x \] such that: \[\Large \frac{du}{dx}=\frac{1}{x} \] might be a bit easier for this kind of integral: \[\Large \int_a^b (\ln x)^2 \frac{1}{x}dx= \int_a^bu^2du \]

OpenStudy (reemii):

"the rest": I rather mean: \(g'\) is the remaining function.